OpenStudy (anonymous):

another question people. . if f(x)= 1/1-x^2 then f(-1/2)equals

OpenStudy (anonymous):

\[f(-\frac{1}{2})=\frac{1}{1-(-\frac{1}{2})^2}\]

OpenStudy (anonymous):

\[=\frac{1}{1-\frac{1}{4}}=\frac{1}{\frac{3}{4}}=\frac{4}{3}\]

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

yeah i got the same answer! thanks(:

OpenStudy (anonymous):

yw